Find the value of: (3/13) ÷ (-4/65)
Find the value of: (3/13) ÷ (-4/65)

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

We have,
= (3/13) × (65/-4) ... [∵reciprocal of (-4/65) is (65/-4)]
The product of two rational numbers = (product of their numerator)/ (product of their denominator)
= (3×65) / (13× (-4))
= 195/-52
= -15/4

Thanks